# Enhancing Brain Age Estimation with a Multimodal 3D CNN Approach Combining Structural MRI and AI-Synthesized Cerebral Blood Volume Data

### v1

@@ 4.1 Performance Improvement with AICBV
Our approach, which integrates predictions from models trained on T1w MRI and AI-generated cerebral blood volume (AICBV), achieved a lower mean absolute error (MAE) than the model proposed by Feng et al. [4]. Their model, which utilized a similar architecture and was trained on the same large and heterogeneous dataset, relied solely on T1w MRI as input. This improved performance underscores not only enhanced predictive accuracy but also the successful incorporation of functional information into BrainAGE estimation—a longstanding objective in the field. AICBV provides an additional independent covariate that a T1w-trained model alone could not replicate due to a lack of prior knowledge of the importance of functional information. By leveraging the DeepC algorithm to generate functional information without requiring additional data collection, our study contributes a novel source of function information to the BrainAGE modeling task.
@@ 4.2 Clinical Implications
The integration of AICBV with structural T1w MRI scans offers significant potential for advancing clinical brain aging assessments. Our model demonstrates that combining functional and structural imaging modalities leads to more accurate predictions of brain age, especially in older populations. This could be particularly valuable in clinical settings for the early detection of neurodegenerative diseases such as Alzheimer’s disease, where subtle functional changes precede visible structural atrophy. Furthermore, the ability to get this functional information from non-invasive MRI scans without the need for further risky measures provides a safer alternative to accessing this valuable context. By providing a more comprehensive view of brain health, this approach has the potential to enhance clinical decision-making, offering a valuable tool for both research and routine clinical practice.
4.3 Clinical Knowledge Alignment
In our linear regression analyses, we systematically evaluated the inclusion and exclusion of specific feature sets: (1) T1w-predicted age only (T-model), (2) AICBV-predicted age only (A-model), (3) combined T1w- and AICBV-predicted ages (TA-model), and (4) combined T1w- and AICBV-predicted ages with sex as an additional covariate (TAS-model). Among these, the TAS-model achieved the lowest MAE of 3.95. Notably, the coefficient for sex in this model was -0.168, with a p-value of 0.02, below the standard significance threshold of 0.05. Given that sex was coded as 0 for female and 1 for male, this negative coefficient suggests that the model adjusts male age predictions downward, indicating that male brains are perceived as appearing older than female brains. This result aligns with established clinical evidence of sex-related differences in brain aging, further validating the model’s anatomical and physiological insights [1, 3].
@@ 5 Conclusion
In this study, we demonstrated that integrating anatomical (T1w MRI) and functional (AICBV) data improves the accuracy of BrainAGE prediction, showcasing the synergistic value of structural and functional data. Additionally, our findings align with clinical understandings of brain aging, further validating the utility of our approach. Future work will explore novel deep learning architectures, including transformer-based models and cross-encoder designs, to more effectively integrate anatomical and functional data and further enhance model performance. These approaches also have the potential to enhance local explainability, providing anatomically grounded insights into predictions. Moreover, the development of more advanced XAI methods, such as utilizing large language models to generate free-text explanations, could make these models more accessible to clinicians, seamlessly integrating explainability into their workflows. Finally, conducting translational studies will be crucial to validate these models in real-world clinical settings and assess their impact on patient care outcomes. By addressing these future directions, we will work to bridge the gap between the cutting-edge machine learning research developments and its practical application in healthcare.

### v3

@@ 4.1 Performance Improvement with AICBV
Our approach, which integrates predictions from models trained on T1w MRI and AICBV, achieved a lower mean absolute error (MAE) than the similar model proposed by Feng et al. [6]. Their model, which utilized a similar architecture and was trained on the same large and heterogeneous dataset, relied solely on T1w MRI as input. This improved performance underscores not only enhanced predictive accuracy but also the successful incorporation of functional information into BrainAGE estimation—a longstanding objective in the field. AICBV provides an additional independent covariate that a T1w-trained model alone could not replicate due to a lack of prior knowledge of the importance of functional information. By leveraging the DeepC algorithm to generate functional information without requiring additional data collection, our study contributes a novel source of function information to BrainAGE modeling.
4.2 Regional Vulnerability and Sex Difference in Brain Aging
In our linear regression analyses, we systematically evaluated the inclusion and exclusion of specific feature sets: (1) T1w-predicted age only (T-model), (2) AICBV-predicted age only (A-model), (3) combined T1w- and AICBV-predicted ages (TA-model), and (4) combined T1w- and AICBV-predicted ages with sex as an additional covariate (TAS-model). Among these, the TAS-model achieved the lowest MAE of 3.95. Notably, the coefficient for sex in this model was -0.168, with a p-value of 0.02, below the standard significance threshold of $\alpha=0.05$. Given that sex was coded as 0 for female and 1 for male, this negative coefficient suggests that the model adjusts male age predictions downward, indicating that male brains are perceived as appearing older than female brains. This result aligns with established clinical evidence of sex-related differences in brain aging, further validating the model’s anatomical and physiological insights [1, 3].
@@ 4.3 Clinical Implications
The integration of AICBV with structural T1w MRI scans offers significant potential for advancing clinical brain aging assessments. Our model demonstrates that combining functional and structural imaging modalities leads to more accurate predictions of brain age, especially in older populations. This could be particularly valuable in clinical settings for the early detection of neurodegenerative diseases such as Alzheimer’s disease, where subtle functional changes precede visible structural atrophy [23]. Furthermore, the ability to get this functional information from non-invasive MRI scans without the need for further risky measures provides a safer alternative to accessing this valuable context. By providing a more comprehensive view of brain health, this approach has the potential to enhance clinical decision-making, offering a valuable tool for both research and routine clinical practice.
BrainAGE has served as a valuable neuroimaging biomarker in clinical practice, offering insights into individual brain aging trajectories [19, 21]. BrainAGE, by estimating the difference between predicted brain age and chronological age, serves as a potential biomarker for detecting neuroanatomical changes that may precede clinical symptoms of neurodegenerative diseases [4]. However, its application faces limitations when only considering structural information in its prediction. Our integrated structural and functional approach captures both the architectural integrity and subtle vascular changes, potentially offering a more sensitive indicator of incipient neurodegenerative trajectories. This integrated biomarker could be invaluable for monitoring disease progression and evaluating the efficacy of novel interventions, guiding clinicians to intervene earlier in the disease course. Our method capitalizes on routine MRI scans and delivers functional-like insights without additional procedures, broadening its potential impact in both preclinical research and everyday patient care.
Additionally, while BrainAGE provides a valuable snapshot of an individual’s current brain health, it does not fully capture the dynamic progression of underlying pathophysiological changes as they unfold over time [7]. Understanding the trajectory of a patient’s aging process—rather than relying on a single timepoint—can be crucial for identifying subtle, cumulative shifts that may signal the early onset or accelerated progression of neurological diseases [14, 22]. By expanding this biomarker to account for longitudinal data, clinicians could tailor its application to individual patients, allowing the model to detect deviations in both structural integrity and functional vascular changes relative to an expected aging paradigm. This personalized, time-sensitive approach has the potential to reveal emerging disease states at an earlier stage, guiding timely intervention and more precise patient management.
@@ 5 Conclusion
In this study, we demonstrated that integrating anatomical (T1w MRI) and functional (AICBV) data improves the accuracy of BrainAGE prediction, showcasing the synergistic value of structural and functional data. Additionally, our findings align with clinical understandings of brain aging, further validating the utility of our approach.
Future work will explore novel deep learning architectures, including transformer-based models and cross-encoder designs, to more effectively integrate anatomical and functional data and further enhance model performance. These approaches also have the potential to enhance local explainability, providing anatomically grounded insights into predictions. Emerging advances in unsupervised and latent space learning can further enhance the clinical applicability of these models. Recent works have leveraged unsupervised deep learning to extract latent vulnerability maps and saliency regions from neuroimaging data [5, 23]. Adapting these latent space approaches to BrainAGE modeling could identify hidden disease signatures and vulnerability pathways without requiring extensive labeled training data. This additional layer of interpretability would not only improve the precision and reliability of brain age predictions, but would also guide clinicians toward more targeted interventions, personalized monitoring regimens, and a deeper understanding of the complex interplay between structural and functional changes in the aging and diseased brain. In this way, the combination of multimodal imaging, robust modeling strategies, and unsupervised learning techniques holds promise for elevating BrainAGE from a research tool into a clinically actionable biomarker.
Moreover, the development of more advanced XAI methods, such as utilizing large language models to generate multimodal, free-text explanations, could make these models more accessible to clinicians, seamlessly integrating explainability into their workflows. Finally, conducting translational studies will be crucial to validate these models in real-world clinical settings and assess their impact on patient care outcomes. By addressing these avenues, we aim to bridge the gap between cutting-edge machine learning research and its practical application in healthcare.
6 Figures
Figure 3: Scatter and residual plots for each model on the test data set.